OpenStudy (chrisplusian):

write the system of equations associated with the matrix given in reduced row echelon form

OpenStudy (chrisplusian):

honestly I don't know how to input this with an equation: Row1: 1,0,0,2 Row2: 0,1,0,3 Riw3: 0,0,1,-2

OpenStudy (anonymous):

System of Equations Standard Form: \[anX + bnY + cnZ = dn\] So: x, y, z, answer Then: Row1: 1,0,0,2 1x + 0y + 0z = 2 Then: Row2: 0,1,0,3 0x + 1y + 0z = 3 Then: Row3: 0,0,1,-2 0x + 0y + 1z = -2 FOLLOWING SO FAR?

OpenStudy (anonymous):

@chrisplusian